OpenStudy (chrisplusian):

If "A" is an nxn matrix show that "A" can be written as the sum of a symmetric matrix and a skew symmetric matrix.

OpenStudy (zzr0ck3r):

True. Do you need to prove it?

OpenStudy (chrisplusian):

yes

OpenStudy (chrisplusian):

I have a definition in my textbook, and it goes like this.....

OpenStudy (chrisplusian):

\[A = V+S, where V= \frac{ 1 }{ 2 }(A + A^T), and S = \frac{ 1 }{ 2 }(A - A^T)\]

OpenStudy (chrisplusian):

I know I can use this but I am wondering if there is another way generically that I can show this is true. I am not supposed to use actual numbers in any proofs, so I can't just do these operations to a matrix and show it works out

OpenStudy (chrisplusian):

I will take a look at the link, thank you